# Gravitational waves from primordial black hole dominance: The effect of inflaton decay rate

### v1

@@ 4 Induced gravitational waves
  • Second-order GWs: Scalar perturbations couple to the tensor ones at second order in perturbation theory and induce GWs. For this reason, these are called scalar-induced gravitational waves (SIGWs). Several mechanisms can produce amplified scalar perturbations [^3] SIGWs are also produced by the perturbations amplified during inflation Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 that collapse into PBH, or even by the evaporation of these PBH Domenech:2024wao ; Domenech:2020ssp .. However, in this work, we focus on the amplified scalar density fluctuations from the PBH-dominated phase and apply the approach in Refs. Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 .
@@
Finally, in Fig. 5, we show the mean mass of the PBH distribution as a function of the fractional density of PBH and the frequency of the peak of the SIGWs produced in each case. To compare, we also show the frequency ranges of some planned GW detectors, such as the Levitated Sensor Detector (LSD) Aggarwal:2020umq , the Einstein Telescope (ET) Maggiore:2019uih , and the Large Interferometer Space Antena (LISA) Amaro-Seoane:2017ekt . This reveals that, for some regions of the parameter space, the frequency of the GWs falls into the detectable range of the LSD and ET detectors. However, their sensitivity is insufficient to detect these GWs, which further motivates their refinement. In addition, we show also the effect of changing the decay rate. A larger decay rate (left plot) implies that PBH have more time to dominate, as the scalar field decays sooner. As a consequence, GWs are produced more abundantly and on a wider span of frequencies. On the contrary, a smaller decay rate (right plot) implies a reduced production of GWs. To produce this plot, we have considered the PBH that have enough time to dominate ($t_{\text{PBH}}<t_{\text{eva}}$), and excluded the scenarios where GWs are overproduced ($\mathcal{I}_{\Omega}^{\text{PBH}}<\mathcal{I}_{\text{BBN}}$). This plot is aimed to be compared with Plot. 3 of Papanikolaou:2020qtd , where the authors study the production of GWs from a PBH-dominated phase without focusing on the production mechanism or the inflationary details. In this work, in addition, we consider the whole evolution of the PBH, from formation mechanism to their domination and final evaporation, along with details of inflaton decay. Since the energy density of SIGWs $I_{\rm GW}$ relies on the PBH domination duration (determined by the inflaton decay rate (through (31)) and the PBH mass fraction (See (13) and (15)), our results are significantly different from Papanikolaou:2020qtd . Therefore, our considerations of PBH formation details and inflaton decay produce high-frequency GWs induced by PBH domination in contrast to the results in Papanikolaou:2020qtd , which indicate a low-frequency domain. Note that the authors in Papanikolaou:2020qtd evaluate the plot at the late matter-radiation equality, whereas Fig. 5 is evaluated at present time.
@@
Figure 5: Mass of the PBH as a function of the frequency at which the fractional energy density of GWs peaks and the initial fractional energy density of PBH. The green, pink, and purple bands correspond to the range of detectable frequencies of the LSD, ET, and LISA GW detectors, respectively.
@@ 5 Conclusions
In this work, we have explored the formation and evolution of PBH in an early matter-dominated universe, focusing on their potential dominance, which provides an alternative reheating mechanism through Hawking radiation. Our approach focuses on the Khlopov-Polnarev formalism Khlopov:1980mg ; Khlopov:1982ef ; Polnarev:1985btg ; Khlopov:2008qy ; Polnarev:1981 ; Carr:2020gox ; Carr:2021bzv ; Harada:2016mhb to describe PBH formation during the preheating (matter-dominated) phase, considering an extended distribution of perturbations rather than a monochromatic one. This scenario differs significantly from the standard ones considered in the literature Papanikolaou:2020qtd ; Domenech:2023jve ; Domenech:2024wao ; Domenech:2020ssp ; Papanikolaou:2022chm , where PBH are considered to form during radiation domination and the mass fraction is computed using the Press-Schechter formalism Press:1973iz . To achieve a significant amount of PBH, we consider a Starobinsky-like inflationary model Starobinsky:1980te with some feature in the potential that amplifies perturbations around a particular scale, parametrized with a Gaussian peak, eqn. (3). PBH domination, in our framework, is affected by the inflaton field decay rate ($\Gamma_{\phi}$) into radiation via a decay rate $\Gamma_{\phi}$, whose values we choose according to the Starobinsky model Jeong:2023zrv . Then, the evolution of the energy densities is solved with a system of coupled Boltzmann equations (24). If PBH dominate for a sufficient amount of time (that depends on inflaton decay rate through (31)), the Poissonian density fluctuations they produce induce GWs at second order in perturbation theory Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 that could reach the BBN bound (47). Our study revealed that the duration of the PBH-dominated phase, and thus the production of GWs phase is highly sensitive to:
@@
We computed the power spectrum of these induced GWs and found that the resulting signal could be within the detectable range of future gravitational wave detectors such as the LSD and ET, see Fig. 5. Further, our results indicate that considering an early matter-dominated phase together with a decay rate for the inflation field allows for a relaxation of earlier constraints on PBH, as Fig. 4 reveals. This suggests that PBH could have played a more significant role in cosmic evolution than early studies indicate, which highlights the importance of considering the interplay between PBH formation during a matter-dominated phase using the Khlopov-Polnarev formalism, the inflaton decay to radiation, and the emission of induced gravitational waves in the early universe.
@@ Appendix B Analytical solutions for $\Omega_{\text{PBH}}(k)$
$B_{1}=A_{1}\left(\frac{8\sqrt{\mathcal{A}_{\text{peak}}}}{5}\right)^{5},$ (56a)
$B_{2}=A_{2}\left(\frac{8\sqrt{\mathcal{A}_{\text{peak}}}}{5}\right)^{6},$ (56b)
$\alpha_{1}=\frac{5}{2\sigma^{2}\ln(10)^{2}},$ (56c)
$\alpha_{2}=\frac{3}{\sigma^{2}\ln(10)^{2}}.$ (56d)

### v2

@@ 4 Induced gravitational waves
  • Second-order GWs: Scalar perturbations couple to the tensor ones at second order in perturbation theory and induce GWs. For this reason, these are called scalar-induced gravitational waves (SIGWs). Several mechanisms can produce amplified scalar perturbations [^4] SIGWs are also produced by the perturbations amplified during inflation Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 that collapse into PBH, or even by the evaporation of these PBH Domenech:2024wao ; Domenech:2020ssp .. However, in this work, we focus on the amplified scalar density fluctuations from the PBH-dominated phase and apply the approach in Refs. Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 ; Inomata:2019ivs .
@@
Figure 5: Mass of the PBH as a function of the frequency at which the fractional energy density of GWs peaks and the initial fractional energy density of PBH. The green, pink, and purple bands correspond to the range of detectable frequencies of the LSD, ET, and LISA GW detectors, respectively. Brighter orange-to-black colors correspond to our results, whereas lighter colors correspond to the results of ref. Papanikolaou:2020qtd .
In Fig. 5, we show the mean mass of the PBH distribution as a function of the fractional density of PBH and the frequency of the peak of the SIGWs produced in each case. To compare, we also show the frequency ranges of some planned GW detectors, such as the Levitated Sensor Detector (LSD) Aggarwal:2020umq , the Einstein Telescope (ET) Maggiore:2019uih , and the Large Interferometer Space Antenna (LISA) Amaro-Seoane:2017ekt . This reveals that, for some regions of the parameter space, the frequency of the GWs falls into the detectable range of the LSD and ET detectors. However, their sensitivity is insufficient to detect these GWs, which further motivates their refinement. In addition, we also show the effect of changing the decay rate. A larger decay rate (left plot) implies that PBH have more time to dominate, as the scalar field decays sooner. As a consequence, GWs are produced more abundantly and on a wider span of frequencies. On the contrary, a smaller decay rate (right plot) implies a reduced production of GWs. To produce this plot, we have considered the PBH that have enough time to dominate ($t_{\text{PBH}}<t_{\text{eva}}$), and excluded the scenarios where GWs are overproduced ($\mathcal{I}_{\Omega}^{\text{PBH}}<\mathcal{I}_{\text{BBN}}$). This plot is aimed to be compared with Plot. 3 of Papanikolaou:2020qtd , where the authors study the production of GWs from a PBH-dominated phase without focusing on the production mechanism or the inflationary details. The results from Fig. 3 of Papanikolaou:2020qtd are shown in lighter colors in Fig. 5. These, in general, correspond to lower frequencies and are placed below our results, shown in brighter colors. In this work, in addition, we consider the whole evolution of the PBH, from formation mechanism to their domination and final evaporation, along with details of inflaton decay. Since the energy density of SIGWs $I_{\rm GW}$ relies on the PBH domination duration (determined by the inflaton decay rate (through (31)) and the PBH mass fraction (See (13) and (15)), our results are significantly different from Papanikolaou:2020qtd . Therefore, our considerations of PBH formation details and inflaton decay produce high-frequency GWs induced by PBH domination in contrast to the results in Papanikolaou:2020qtd , which indicate a low-frequency domain. Note that the authors in Papanikolaou:2020qtd evaluate the plot at the late matter-radiation equality, whereas Fig. 5 is evaluated at the present time.
Finally, and for completeness, we show in Fig. 6 the GW spectrum as a function of the frequency for the case of a decay rate of $\Gamma_{\phi}=10^{-19}M_{\text{pl}}$ and a mean PBH mass of $M_{\text{PBH},0}=10^{6}$g. The result is shown for several values of $\mathcal{A}_{\text{peak}}$, which translates into different mass fractions of PBH, $\Omega_{\text{PBH}}(t_{\text{in}})$. As explained above, the GW spectrum lies below the sensitivity of the future GW detectors. However, it is worth mentioning that one could also consider detectors based on electromagnetic resonant cavities Herman:2022fau . Several such detectors, both operational and planned, have been proposed Gatti:2024mde ; Schneemann:2024qli , offering very high experimental sensitivities. Nevertheless, their characteristic frequency ranges lie in the GHz region of the spectrum, slightly above the upper limit of the predicted signal in our scenario.
Figure 6: An example of GW spectra for the scenario where $\Gamma_{\phi}=10^{-19}M_{\text{pl}}$ and $M_{\text{PBH},0}=10^{6}$g shown for several values of the amplitude of the Gaussian peak $\mathcal{A}_{\text{peak}}$.
@@ 5 Conclusions
In this work, we have explored the formation and evolution of PBH in an early matter-dominated universe, focusing on their potential dominance, which provides an alternative reheating mechanism through Hawking radiation. Our approach focuses on the Khlopov-Polnarev formalism Khlopov:1980mg ; Khlopov:1982ef ; Polnarev:1985btg ; Khlopov:2008qy ; Polnarev:1981 ; Carr:2020gox ; Carr:2021bzv ; Harada:2016mhb to describe PBH formation during the preheating (matter-dominated) phase, considering an extended distribution of perturbations rather than a monochromatic one. This scenario differs significantly from the standard ones considered in the literature Papanikolaou:2020qtd ; Domenech:2023jve ; Domenech:2024wao ; Domenech:2020ssp ; Papanikolaou:2022chm ; Papanikolaou:2024kjb ; He:2024luf , where PBH are considered to form during radiation domination and the mass fraction is computed using the Press-Schechter formalism Press:1973iz . To achieve a significant amount of PBH, we consider a Starobinsky-like inflationary model Starobinsky:1980te with some feature in the potential that amplifies perturbations around a particular scale, parametrized with a Gaussian peak, eqn. (3). PBH domination, in our framework, is affected by the inflaton field decay into radiation via a decay rate $\Gamma_{\phi}$, whose values we choose according to the Starobinsky model Jeong:2023zrv . Then, the evolution of the energy densities is solved with a system of coupled Boltzmann equations (24). If PBH dominate for a sufficient amount of time (that depends on inflaton decay rate through (31)), the Poissonian density fluctuations they produce induce GWs at second order in perturbation theory Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 that could reach the BBN bound (47). Our study revealed that the duration of the PBH-dominated phase, and thus the production of GWs phase is highly sensitive to:
@@
We computed the fractional energy density of these induced GWs and found that the resulting signal could be within the detectable frequency range of future gravitational wave detectors such as the LSD and ET, see Fig. 5. However, as shown in Fig. 6, the experimental sensitivity of these detectors is insufficient to detect the computed signal, highlighting the need for further improvements or the design of new instruments in this aspect. Further, our results indicate that considering an early matter-dominated phase together with a decay rate for the inflation field allows for a relaxation of earlier constraints on PBH, as Fig. 4 reveals. This suggests that PBH could have played a more significant role in cosmic evolution than early studies indicate, which highlights the importance of considering the interplay between PBH formation during a matter-dominated phase using the Khlopov-Polnarev formalism, the inflaton decay to radiation, and the emission of induced gravitational waves in the early universe.
@@ Appendix B Analytical solutions for $\Omega_{\text{PBH}}(k)$
$B_{1}=A_{1}\left(\frac{8\sqrt{\mathcal{A}_{\text{peak}}}}{5}\right)^{5},\quad B_{2}=A_{2}\left(\frac{8\sqrt{\mathcal{A}_{\text{peak}}}}{5}\right)^{6}$ (56a)
$\alpha_{1}=\frac{5}{2\sigma^{2}\ln(10)^{2}},\quad\alpha_{2}=\frac{3}{\sigma^{2}\ln(10)^{2}}.$ (56b)